OpenStudy (anonymous):

272=16n can i get any help here please and thanks

OpenStudy (anonymous):

Divide 272 by 16 to get n alone

OpenStudy (anonymous):

i did that but like, i got -256 and i divided after that by sixteen and now i'm lost

OpenStudy (tkhunny):

If you divide, the sign will not change. 272/16 = 136/8 = 68/4 = 34/2 = 17

OpenStudy (anonymous):

you must've subtracted it

OpenStudy (anonymous):

OH WAIT OKAY I SEE NOW

OpenStudy (anonymous):

i'm still partially confused about tkhunny's reply though. can you explain a little more in depth?

OpenStudy (tkhunny):

272/16 was good enough. I was just simplifying the fraction. You should be able to do that. Or, just use a calculator and see what you get.

OpenStudy (anonymous):

So you basically have to get n by itself. so that means that you would have to divide 16, on each side to even out the equation, THEN you would get n alone and have your answer. so 272/16= 17

OpenStudy (anonymous):

n=17.

OpenStudy (anonymous):

oh okay i see now sweet thanks man